OpenStudy (anonymous):

What happens, though, if you are dealing with a system that has either really weird numbers or really larger numbers? How could you use the other two methods to supplement the use of graphing to find the necessary solution to the system?

OpenStudy (anonymous):

I was trying to post earlier, but the site wouldn't let my message go through (if it did, I still don't see it posted). I was going to ask if you could post the previous information that you had to make this question. I'm not sure what the "other two methods to supplement the use..." are.

OpenStudy (anonymous):

Graphing is definitely a great way to visualize systems of equations and to get a good idea about the nature of the solution of the system. I think that given the nature of today's technology (i.e. graphing calculators, computers, etc.), graphing techniques can be improved to have higher levels of accuracy. What happens, though, if you are dealing with a system that has either really weird numbers or really larger numbers? How could you use the other two methods to supplement the use of graphing to find the necessary solution to the system?

OpenStudy (anonymous):

I'm afraid that we're still missing what this sentence is referring to: "How could you use the other two methods to supplement the use of graphing "

OpenStudy (anonymous):

elimination method is one

OpenStudy (anonymous):

Ah, ok. Elimination and substitution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome. Alright. Let's say that the answer you want on a graph is on somewhere like (3, 3.7416573). Unfortunately, we can't tell straight from the graph what the answer is.

OpenStudy (anonymous):

The reason you'd want to use substitution/elimination is get something like this solved to an exact answer:

OpenStudy (anonymous):

2*sqrt(14)+4 + x = 2y x + y = 0 y = 0*5 By using substitution, using equation 3, you will know that y = 0, so equation 2 becomes 0 + y = 0

OpenStudy (anonymous):

Ah crud... bad example.

OpenStudy (anonymous):

Basically...... the reason you can't always rely on graphs is because you won't always get pretty answers like (4,0). Sometimes you might get (4,3.7416573), which even most graphing calculators don't know is really "Square root of 14". But with elimination and substitution, you can get the exact answer.